A plane flies 900 miles in 3 hours with a tail wind. It takes the sme plane 5 hours to fly 900 miles against the wind. Find the rate of the plane in calm air and the rate of the wind.
Answer by Paul(988)

x = rate of the plane
y = rate of the wind
3(x+y)=900
5(x-y)=900
3x+3y=900----------------------->5(3x+3y)=900 Find the same number to eliminate
5x-5y=900----------------------->3(5x-5y)=900

15x+15y=4500
15x-15y=2700
30x=7200
x=240

5(240-y)=900
1200-5y=900
-5y=-300
y=60

Hence, the speed of the plane is 240 miles per hour and the rate of the wind is 60 miles per hour.